# An Introduction to Reinforcement Learning: Fundamental Concepts and Practical Applications

### v1

@@ Abstract
Reinforcement Learning (RL) is a branch of Artificial Intelligence (AI) which focuses on training agents to make decisions by interacting with their environment to maximize cumulative rewards. An overview of RL is provided in this paper, which discusses its core concepts, methodologies, recent trends, and resources for learning. We provide a detailed explanation of key components of RL such as states, actions, policies, and reward signals so that the reader can build a foundational understanding. The paper also provides examples of various RL algorithms, including model-free and model-based methods. In addition, RL algorithms are introduced and resources for learning and implementing them are provided, such as books, courses, and online communities. This paper demystifies a comprehensive yet simple introduction for beginners by offering a structured and clear pathway for acquiring and implementing real-time techniques.
@@ 1 Introduction
Reinforcement Learning (RL) is another type of learning in AI that focuses on training by interacting with the environment. In contrast to supervised learning, where an agent learns from labeled examples, or unsupervised learning, which is based on detecting patterns in the data, RL deals with an autonomous agent that must make intuitive decisions and consequently learn from its actions. The key idea is to learn how the world works (e.g., what action gets a reward and which does not) to maximize cumulative rewards over time through trial-and-error exploration [sutton2018reinforcement]. This paper will serve as a crash course in RL, with the aim that readers should be geared up for solving real-world problems using it. This paper considers the readers to have basic knowledge of Machine Learning (ML) algorithms like Supervised and Unsupervised Learning.
RL revolves around several key concepts: States, Actions, Policies, and Rewards. Each of these components plays a crucial role in defining the agent’s interaction with its environment and the overall learning process.
@@
A State (s) represents a specific condition or configuration of the environment at a given time as perceived by the agent. A state sets the scene for the agent to make choices and select actions. This is the state space, which describes all states from which an agent can choose for each action that occurs. For example, in chess, a state might be one specific layout of pieces on the board. Actions (a) are the set of possible moves or decisions an agent can make while interacting with the environment. A selected action is part of the strategy followed by an agent to reach its desired goals according to its current states and policy. In the chess example, moving a piece from one square to another is an action. A policy ($\pi$) guides the behavior of a learning agent by mapping perceived states of the environment into actions. This could be a simple function, a lookup table, or a complex computation. Policies can also be stochastic, defining the likelihood of taking certain actions. A deterministic policy assigns a specific action to each state, while a stochastic policy assigns probabilities for all potential actions in any given state. The Rewards (r) are a critical factor in RL because it provides the agent with an objective at each time step, defining both local and global goals that the agent aims to achieve over time. Rewards differentiate positive from negative events and help update policies according to the outcomes of actions. Rewards depend on the state of the environment and the actions taken. For instance, in a game scenario, winning points might be the reward, while losing points could be the penalty. Environment Model is an approximation or prediction of the environment, mapping out what will be available (the next state and reward) given a particular input from the state and action. These models help with planning by identifying what actions should be taken based on possible future events. RL approaches that use models are called model-based methods, whereas those relying solely on trial-and-error learning are model-free methods.
To cement these concepts, let us consider an RL agent navigating a maze. The state represents the agent’s current position within the maze. The actions could be moving north, south, east, or west. The policy dictates which direction the agent should move based on its current state. The reward signal might give a positive reward when the agent reaches the end of the maze and negative rewards for hitting walls or making incorrect move.
@@
Figure 1 illustrates the interaction between an RL agent and its environment. The agent observes the current state ($S_{t}$), selects an action ($A_{t}$) based on its policy, and receives a reward ($R_{t}$) along with the next state ($S_{t+1}$). This feedback loop is critical for the agent to learn and update its policy to maximize cumulative rewards.
Figure 1: Overview model of Reinforcement Learning, influenced by [sutton2018reinforcement]
Several critical challenges must be addressed in order to understand and implement RL. A key challenge is balancing exploration with exploitation. As part of exploration, the agent tries new actions in order to determine their effects and gather more information about the environment. To learn and improve the agent’s understanding of the potential rewards of different actions, this information is essential. Alternatively, exploitative behavior involves the agent using its current knowledge to make decisions that maximize immediate rewards. In order to achieve the highest possible reward, the agent utilizes the best-known actions in accordance with its learned policy [thrun1992efficient]. Traditionally, the Epsilon-Greedy method is used to resolve this trade-off. An agent chooses a random action with a probability of $\epsilon$ (exploration) and the best-known action with a probability of $1-\epsilon$ (exploitation). The value of $\epsilon$ may be fixed or may decay over time [arulkumaran2017brief, sutton2018reinforcement].
To solve problems with RL, we have to formulate them by Markov decision processes (MDPs). Using an MDP, one can model decision-making where outcomes are partly determined by chance and partly controlled by the decision-maker. As a result, the core idea is to capture the fundamental challenges faced by an agent interacting with its environment in order to achieve its goals. It is essential that the agent perceives the environment’s state, takes actions which affect it, and has objectives related to the environment’s states that are embodied in the reward signal. The process of forming a problem in MDP often poses a variety of challenges, which can be viewed as the second challenge[sutton2018reinforcement].
The rest of the paper is organized as follows: Section 2 introduces multi-bandits, that are helpful in understanding the nature of RL problems by starting with an easy example. A thorough introduction to Markov Decision Processes is given in section 3, a fundamental step in learning RL. A complete introduction to policies and value functions, core ideas of every RL algorithms, is given in section 4. Section 5 examines the limitation of policy iteration by introducing a new concept called Value Iteration. Understanding various methods in RL is pivotal to learning and implementing RL. Hence, section 6 gives complete descriptions about each method. In section 7, we outline the advancements in RL from its early foundations to recent trends, giving readers a comprehensive understanding of the field’s development. Additionally, Also, must-know widely-used algorithms are provided in section 8. Finally, useful resources for learning RL are provided to avoid perplexion.
The following section explore simple problems involving finite states and actions, followed by a discussion of the main problem framework, finite Markov decision processes. This includes key concepts like Bellman equations and value functions essential for decision-making problems.
2 K-Armed Bandits
In this section, we examine a simple RL form with a single state called bandit problems to broaden our understanding of RL in a simple environment to cement the mentioned points in the former section.
Firstly, we need to understand what is the difference between types of feedback, as the reward signal is a form of a feedback received by agents. In Figure 2, an agent interacts with the environment, takes observations ($O_{t}$) at each timestep, and based on these observations, performs actions. There are four possible actions: $A_{1},A_{2},A_{3},$ and $A_{4}$. Assume the best action $a^{*}$ is $A_{2}$, but the agent chose $A_{4}$.
Figure 2: Agent interaction in a simple environment of four-armed bandit
Instructive feedback indicates the chosen action is wrong and the best action is $A_{2}$, typical in supervised learning tasks. Evaluative feedback only indicates the chosen action is not good, typical in RL tasks, and evaluative feedback depends entirely on the actions taken. The stated example shows the difference of the feedback received in RL environment.
In the following sections, evaluating action-values will be examined in a simple form of RL, which is Immediate Reinforcement Learning (IRL).
In order to understand the K-armed bandit problem or IRL, assume that you are required to select from a number of options (actions) repeatedly. You receive a numerical reward (or punishment) after each choice based on a stationary probability distribution. The concept of ”stationarity” refers to the fact that probability distributions for rewards and transitions remain unchanged over time. For the purpose of simplifying learning, RL algorithms rely on stationarity and model-free methods such as Q-learning. However, this assumption does not always hold in real-world applications, so algorithms for non-stationary environments are required [li2022testing].
For each of the $k$ actions available, there is an expected average reward, referred to as the value of the action [sutton2018reinforcement]. Let us denote the action chosen at time step $t$ as $A_{t}$ and the reward received as $R_{t}$. Define $q^{*}(a)$ as the value of an arbitrary action $a$.
The value of an action $q^{*}(a)$ is defined as the expected reward when the agent takes action $a$ at any given time step $t$. Mathematically, this is represented as:
$q^{*}(a)=\mathbb{E}[R_{t}\mid A_{t}=a]$ (1)
Here, we are defining $q^{*}(a)$ as the expected reward for taking action $a$. This forms the basis for evaluating actions in the k-armed bandit problem.
If we knew the best action to take, solving the problem would be trivial as we would always choose the best action. Without this knowledge, we must estimate each action’s value, denoted as $Q_{t}(a)$ at time step $t$, which should approximate $q^{*}(a)$.
@@ 2 K-Armed Bandits
After estimating action values, there is at least one action with the highest estimated value at each time step. These are called greedy actions. Selecting a greedy action exploits current knowledge for immediate reward, while selecting non-greedy actions explores to improve estimates. Balancing exploration and exploitation is crucial in RL. As mentioned earlier, opting for exploitation maximizes immediate reward, while exploration can yield higher overall reward depending on various factors [sutton2018reinforcement, thrun1992efficient].
Estimating action values through sampling averages is efficient for stationary bandit problems. In real-world RL problems with non-stationary environments, it makes sense to weight recent rewards more. Using a constant step-size parameter is popular. Rewriting the sample-average equation as:
To update the action value, we use the following update rule:
$Q_{n+1}\equiv Q_{n}+\alpha[R_{n}-Q_{n}]$ (2)
Here:
  • $Q_{n+1}$: The updated action value after the $n$-th reward.
  • $Q_{n}$: The current action value before the $n$-th reward.
  • $\alpha$: The learning rate, a parameter that determines the extent to which new information overrides the old information.
  • $R_{n}$: The reward received after taking an action at step $n$.
This update rule incrementally adjusts the action value based on the difference between the received reward and the current estimate, weighted by the learning rate $\alpha$. This iterative process allows the agent to refine its value estimates and improve decision-making over time.
To estimate action-values in K-armed bandit problems, we use the following equation, which represents the action value $Q_{t}(a)$ as the average of the rewards received from that action up to time step $t$.
We explore techniques for assessing action value and making choices based on these estimates, known as action-value methods. The true value of an action corresponds to its average reward upon selection, estimated by averaging received rewards, as written in equation below:
$Q_{t}(a)=\frac{\texttt{sum of rewards when the action }a\texttt{ is taken prior to time step }t}{\texttt{number of times the action }a\texttt{ is taken prior to time step }t}$ (3)
$Q_{t}(a)$ can be mathematically expressed as:
$Q_{t}(a)=\frac{\sum_{i=1}^{t-1}R_{i}\cdot\mathbf{1}_{\{A_{i}=a\}}}{\sum_{i=1}^{t-1}\mathbf{1}_{\{A_{i}=a\}}}$ (4)
where $\mathbf{1}_{\{A_{i}=a\}}$ is the indicator function. The indicator function $\mathbf{1}_{\{A_{i}=a\}}$ is used to count the number of times action $a$ has been taken up to time step $t$.
@@
$\mathbf{1}_{\{A_{i}=a\}}=\begin{cases}1&\texttt{if action }a\texttt{ is taken at time step }i\\ 0&\texttt{otherwise}\end{cases}$ (5)
If the denominator is zero, $Q_{t}(a)$ is defined as a default value (e.g., 0). It is noted to be that as the denominator approaches infinity, $Q_{t}(a)$ converges to $q^{*}(a)$ by the law of large numbers [erd1970new].
An estimated action value is calculated based on an average of reward samples using the sample-average method. This method may not be the most effective, but it can be used as a starting point for selecting actions based on the estimates [sutton2018reinforcement].
$A_{t}\equiv\arg\max_{a}Q_{t}(a)$ (6)
Here, $A_{t}$ represents the action selected at time step $t$, which is the action with the highest estimated value $Q_{t}(a)$.
Greedy selection exploits current knowledge to maximize immediate reward. However, it can be problematic as it requires exploring all actions to know their rewards. A simple yet efficient alternative is being greedy most of the time but selecting randomly from all actions with a small probability, $\epsilon$. This method, called the $\epsilon$-greedy method, balances exploration and exploitation.
These methods ensure all actions are sampled enough to estimate their true value accurately, making the selection of the best action almost certain over time. However, these are theoretical long-term benefits and may not directly indicate practical effectiveness [ladosz2022exploration].
Estimating action values by sampling averages is inefficient for large numbers of samples ($K>n$). A more efficient approach derives $Q_{n}$ as:
Now, the action value using the following formula for the sample average is defined in a following way:
$Q_{n}\equiv\frac{R_{1}+R_{2}+\cdots+R_{n-1}}{n-1}$ (7)
We can express the update rule for action value recursively as follows:
$Q_{n+1}=Q_{n}+\frac{1}{n}[R_{n}-Q_{n}]$ (8)
This recursive equation requires memory only for $Q_{n}$ and $n$, with minimal computation for each reward. The general form of the update rule for action value is:
@@
The target suggests a preferred direction for adjustment, despite noise. In the example, the target is the $n$th reward [luo2024survey].
@@
Sample averages apply to stationary bandit problems. In non-stationary environments, recent rewards are more relevant. Using a constant step-size parameter, the equation can be rewritten as:
$Q_{n+1}\equiv Q_{n}+\alpha[R_{n}-Q_{n}]$ (10)
@@
The initial estimates of action values, $Q_{1}(a)$, play a crucial role in the learning process. These initial guesses influence the early decisions made by the agent. While sample-average methods can reduce this initial bias after each action is chosen at least once, methods using a constant step-size parameter, $\alpha$, tend to mitigate this bias more gradually over time. Setting optimistic initial values can be particularly advantageous. By assigning higher initial estimates (e.g., +5), the agent is encouraged to explore more actions early on. This is because the initial optimism makes untried actions appear more attractive, thus promoting exploration even when the agent uses a greedy strategy. This approach helps ensure that the agent thoroughly investigates the action space before converging to a final policy.
However, this strategy requires careful consideration in defining the initial values, which are often set to zero in standard practice. The choice of initial values should reflect an informed guess about the potential rewards, and overly optimistic values can prevent the agent from converging efficiently if not properly managed.
Overall, optimistic initial values can be a useful technique to balance exploration and exploitation in RL, encouraging broader exploration and potentially leading to more optimal long-term policies [wiering2012reinforcement].
As a result of the uncertainty associated with the estimation of action values, exploration is essential. In the $/epsilon$-greedy method, non-greedy actions are explored indiscriminately. It is preferable to explore non-greedy actions selectively based on their potential optimality and uncertainty. Based on the Upper Confidence Bound (UCB) method, actions are selected based on the following criteria:
$A_{t}\equiv\arg\max_{a}\left[Q_{t}(a)+c\sqrt{\frac{\ln t}{N_{t}(a)}}\right]$ (11)
Here, $\ln t$ denotes the natural logarithm of $t$, $c>0$ controls exploration, and $N_{t}(a)$ is the number of times action $a$ has been taken prior to time step $t$. If $N_{t}(a)=0$, $a$ is considered a maximizing action.
As part of its value estimation, UCB incorporates uncertainty into its calculation of an action’s upper limit. Confidence levels are adjusted by the constant $c$. By choosing action $a$, the uncertainty associated with $a$ decreases as $N_{t}(a)$ increases, whereas by choosing other actions, the uncertainty associated with $a$ increases as $t$ increases. Using the natural logarithm, all actions are eventually explored as uncertainty adjustments decrease over time. The frequency with which actions with low estimates or frequent selections are chosen decreases over time [garivier2011upper].
3 Finite Markov Decision Process

### v2

@@ Abstract
Reinforcement Learning (RL), a subfield of Artificial Intelligence (AI), focuses on training agents to make decisions by interacting with their environment to maximize cumulative rewards. This paper provides an overview of RL, covering its core concepts, methodologies, and resources for further learning. It offers a thorough explanation of fundamental components such as states, actions, policies, and reward signals, ensuring readers develop a solid foundational understanding. Additionally, the paper presents a variety of RL algorithms, categorized based on the key factors such as model-free, model-based, value-based, policy-based, and other key factors. Resources for learning and implementing RL, such as books, courses, and online communities are also provided. By offering a clear, structured introduction, this paper aims to simplify the complexities of RL for beginners, providing a straightforward pathway to understanding and applying real-time techniques.
@@ 1 Introduction
Reinforcement Learning (RL) is a subfield of Artificial Intelligence (AI) that focuses on training by interacting with the environment, aiming to maximize cumulative reward over time [1]. In contrast to supervised learning, where the objective is to learn from labeled examples, or unsupervised learning, which is based on detecting patterns in the data, RL deals with an autonomous agent that must make intuitive decisions and consequently learn from its actions, often without existing data. The key idea is to learn how the world works (e.g., what action gets a (positive) reward and which does not) to maximize cumulative rewards over time through trial-and-error exploration. RL revolves around several key concepts: States, Actions, Policy, Rewards, Transition Dynamics(Probabilities), and Environment Model. Each of these components plays a crucial role in defining the agent’s interaction with its environment and the overall learning process, and are defined over the next paragraphs. This paper assumes readers to have basic knowledge of Machine Learning (ML) algorithms like Supervised and Unsupervised Learning.
@@
A State ($s\in\mathcal{S}$) represents a specific condition or configuration of the environment at a given time as perceived by the agent. A state sets the scene for the agent to make choices and select actions, which describes all states from which an agent can choose for each action that occurs. For example, in chess, a state might be one specific layout of pieces on the board. In brief, states are the situations that an agent can be in and observe to make decisions. States can be discrete (the given example) and continuous (e.g., the position of a robot, in terms of $x$ and $y$ coordinates). Actions ($a\in\mathcal{A}$) are the set of possible moves or decisions an agent can make while interacting with the environment. A selected action is part of the strategy followed by an agent to reach its desired goals according to its current states and policy. In the chess example, moving a piece from one square to another is an action. Similar to states, actions can be discrete and continuous. The given chess example has a set of discrete actions (moving the piece according to the rules, which has a finite set of actions). For continuous actions, let us consider the robot example. A robot can change its coordinates and its movement speed, which both are continuous. Policy ($\pi$) guides the behavior of a learning agent by mapping perceived states of the environment into actions. This could be a simple function, a lookup table, or a complex computation, e.g., approximated by Neural Networks. Policies can be stochastic, defining the likelihood of taking certain actions. A stochastic policy ($\pi(a|s)$) defines a probability distribution over possible actions for a given state. Instead of selecting a single action, it samples actions based on the probabilities defined by the policy. On the other hand, A deterministic policy ($\pi(s)=a$) directly maps a state to a specific action. For a given state, the action taken is fixed and does not involve any randomness. In the game of chess, an agent’s policy determines the move it makes given the current board configuration (state). Imagine an agent evaluating its options in a given state of the board. It assigns probabilities to each legal move based on their estimated utility. For instance: $\pi(a_{1}|s)=0.5$, $\pi(a_{2}|s)=0.3$, $\pi(a_{3}|s)=0.2$. Here, $a_{1}$, $a_{2}$, and $a_{3}$ are possible moves, and the policy specifies a 50% chance of selecting $a_{1}$, 30% for $a_{2}$, and 20% for $a_{3}$. The agent chooses its move by sampling from this distribution. This approach introduces diversity in decision-making, which can be particularly useful when exploring novel strategies or dealing with uncertain opponents. Conversely, a deterministic policy maps the current state $s$ to a specific move $a$ (e.g., $\pi(s)=a_{1}$). In this case, the agent always chooses $a_{1}$ (e.g., moving the queen to a specific square) when the board configuration matches $s$. Deterministic policies are efficient in well-defined and predictable scenarios, such as when the agent has already learned the optimal moves for most board configurations. Rewards ($r\in\mathcal{R}$) are a critical factor in RL since it provides the agent with an objective at each time step, defining both local and global goals that the agent aims to achieve over time. Rewards differentiate positive from negative events and help update policies according to the outcomes of actions. Rewards depend on the state of the environment and the actions taken. For instance, in a Chess scenario, winning the game will lead to a positive reward, while losing points will lead to a negative reward. Also, the game might end up in a draw, which lead to no reward. Reward function defines the immediate reward obtained when taking action $a$ in state $s$. It can be deterministic ($\mathcal{R}(s,r)=r$) or stochastic ($P(r|s,a)$). Transition Dynamics function defines the probability of reaching a new state $s^{\prime}$ given the current state $s$ and action $a$. It can be written mathematically as $\mathcal{P}(s^{\prime}|s,a)$. Environment Model is an approximation or prediction of the environment, mapping out what will be available (the next state and reward) given a particular input from the state and action. These models help with planning by identifying what actions should be taken based on possible future events. RL approaches that use models are called model-based methods, whereas those relying solely on trial-and-error learning are model-free methods. This characteristic is an important factor to choose the appropriate algorithms, suitable for the problem. The environment’s dynamics can be represented as a transition function and a reward function. Together, these form the Markov Decision Process (MDP):
$\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$
where $\mathcal{S}$ is state space, $\mathcal{A}$ is action space, $\mathcal{P}(s^{\prime}|s,a)$ is the transition dynamics, $\mathcal{R}(s,a)$ is the Reward function, and $\gamma\in\left[0,1\right]$. $\gamma$ is the discount factor, a parameter that determines how much importance is given to future rewards compared to immediate rewards.
@@
Figure 1 illustrates the interaction between an RL agent and its environment. The agent observes the current state ($\mathcal{S}_{t}$), selects an action ($\mathcal{A}_{t}$) based on its policy ($\pi$), and receives a reward based on the reward distribution ($\mathcal{R}_{t}$) along with the next state ($\mathcal{S}_{t+1}$). This feedback loop is critical for the agent to learn and update its policy to maximize cumulative rewards.
Figure 1: Overview of Reinforcement Learning [1]
The rest of the paper is organized as follows: Section 2 introduces backgrounds and ket concepts in RL, starting with multi-armed bandit. Bandits are a great place to start learning RL’s foundational concepts like value functions and the Bellman equations. A complete introduction to core RL methods is given in section 3. Section 4 analyzes essential RL algorithms, categorized them in a comprehensive manner. In section 5, useful resources for learning RL are provided for readers would like to delve deeper in the realm of RL. Finally, section 6 Concludes the paper.
2 Backgrounds & Key Concepts
Understanding RL requires a solid grasp of its foundational principles and the mathematical frameworks that guide agent-environment interactions. This section introduces the essential building blocks of it, starting with the simplest form of decision-making under uncertainty: the Multi-Armed Bandit problem. This provides an intuitive gateway to understanding how agents learn from evaluative feedback. Building on this, we delve into MDPs, which formalize sequential decision-making by balancing immediate and future rewards. Finally, we discuss key RL metrics, such as value functions and policies, which form the backbone of most RL algorithms. Together, these topics lay the groundwork for exploring the core RL methods in the next section.
2.1 Multi-Armed Bandit(s)
The Multi-Armed Bandit problem serves as a foundational example for understanding Reinforcement Learning (RL). It represents a simplified decision-making scenario where an agent repeatedly chooses from $\mathcal{K}$ actions (or arms) to maximize cumulative rewards over time. Each action is associated with an unknown reward distribution, requiring the agent to balance exploration (gathering information about all actions) and exploitation (maximizing immediate rewards using the best-known action). Unlike supervised learning, where instructive feedback explicitly points to correct actions, RL often involves evaluative feedback. This type of feedback only signals the quality of chosen actions, making it harder to identify optimal strategies without extensive interaction with the environment.
In the $\mathcal{K}$-armed bandit problem, each action $a\in\{a_{1},a_{2},\dots,a_{k}\}$ yields a reward drawn from a stationary probability distribution. The stationarity assumption implies that the reward probabilities remain constant over time [2]. For each of the $k\in\mathcal{K}$ actions available, there is an expected average reward, referred to as the value of the action [1]. Let $\mathcal{A}_{t}$ denote the action chosen at time step $t$, and $\mathcal{R}_{t}$ the corresponding reward received. The true value of an action, $q_{*}(a)$, is defined as the expected reward by taking that action as:
$q_{*}(a)=\mathbb{E}[\mathcal{R}_{t}\mid\mathcal{A}_{t}=a]$ (1)
@@ 2.1 Multi-Armed Bandit(s)
If $q_{*}(a)$ were known for all $a$, the problem would be trivial: the agent would always choose the action with the highest $q_{*}(a)$. However, in practice, these values are unknown and must be estimated through interaction with the environment. After estimating action values, there is at least one action with the highest estimated value at each time step. These are called greedy actions. Selecting a greedy action exploits current knowledge for immediate reward, while selecting non-greedy actions explores to improve estimates. Balancing exploration and exploitation is crucial in RL. Opting for exploitation maximizes immediate reward, while exploration can yield higher overall reward depending on various factors [1, 3]. Estimating action values through sampling averages is efficient for stationary bandit problems. In real-world problems with non-stationary environments, it makes sense to weight recent rewards more. Using a constant step-size parameter is popular. The following update rule can be used to update the action value, which is known as sample-average method.
$\mathcal{Q}_{t}(a)=\frac{\sum_{i=1}^{t-1}\mathcal{R}_{i}\cdot\mathbf{1}_{\{\mathcal{A}_{i}=a\}}}{\sum_{i=1}^{t-1}\mathbf{1}_{\{\mathcal{A}_{i}=a\}}},$ (2)
where $\mathbf{1}_{\{\mathcal{A}_{i}=a\}}$ is an indicator function that equals 1 if action $a$ was chosen at step $i$, and 0 otherwise. Over time, $\mathcal{Q}_{t}(a)$ converges to $q_{*}(a)$ by the law of large numbers [4], provided all actions are sampled sufficiently often.
@@
$\mathbf{1}_{\{\mathcal{A}_{i}=a\}}=\begin{cases}1&\texttt{if action }a\texttt{is taken at time step }i\\ 0&\texttt{otherwise}\end{cases}$ (3)
Equation 2 can be written in a different way as follows, which represents the action value $\mathcal{Q}_{t}(a)$ as the average of the rewards received from that action up to time step $t$.
$\mathcal{Q}_{t}(a)=\frac{\texttt{sum of rewards when the action }a\texttt{ is taken prior to time step }t}{\texttt{number of times the action }a\texttt{ is taken prior to time step }t}$ (4)
$\mathcal{Q}_{n+1}=\mathcal{Q}_{n}+\alpha[\mathcal{R}_{n}-\mathcal{Q}_{n}],$ (5)
@@
One of the critical challenges in solving the multi-armed bandit problem, and in RL as well, is balancing exploration and exploitation, the greedy action selection ($\mathcal{A}_{t}=\arg\max_{a}\mathcal{Q}_{t}(a)$). Several strategies address this trade-off. We touch some of the widely used ones here, starting with $\epsilon-$greedy. In this method, with probability $\epsilon$, choose a random action (exploration); otherwise, select the greedy action (exploitation). This ensures that all actions are sampled, albeit less frequently for suboptimal actions. Another method, known as Upper Confidence Bound (UCB), which balances exploitation and exploration by incorporating uncertainty into action selection [5]:
$\mathcal{A}_{t}=\arg\max_{a}\left[\mathcal{Q}_{t}(a)+c\sqrt{\frac{\ln t}{\mathcal{N}_{t}(a)}}\right],$ (6)
where $\mathcal{N}_{t}(a)$ is the number of times action $a$ has been chosen, and $c>0$ controls the exploration rate. UCB prioritizes actions with high uncertainty or fewer samples. Optimistic Initialization, assigns high initial values to $\mathcal{Q}_{1}(a)$, encouraging exploration of all actions early on. For example, setting $\mathcal{Q}_{1}(a)=+5$ makes untried actions appear attractive initially.
Algorithm 1 A simple bandit algorithm
1:  Initialize:
2:  for $a=1$ to $k$ do
3:     $\mathcal{Q}(a)\leftarrow 0$
4:     $\mathcal{N}(a)\leftarrow 0$
5:  end for
6:  Repeat forever:
7:  $\mathcal{A}\leftarrow\begin{cases}\arg\max_{a}\mathcal{Q}(a)&\text{with probability }1-\epsilon\\ \text{a random action}&\text{with probability }\epsilon\end{cases}$
8:  $R\leftarrow\texttt{bandit}(\mathcal{A})$
9:  $\mathcal{N}(\mathcal{A})\leftarrow\mathcal{N}(\mathcal{A})+1$
10:  $\mathcal{Q}(\mathcal{A})\leftarrow\mathcal{Q}(\mathcal{A})+\frac{1}{\mathcal{N}(\mathcal{A})}\big{[}\mathcal{R}-\mathcal{Q}(\mathcal{A})\big{]}$
These methods ensure all actions are sampled enough to estimate their true value accurately, making the selection of the best action almost certain over time. However, these are theoretical long-term benefits and may not directly indicate practical effectiveness [6]. To begin with, estimating action values by sampling averages is inefficient for large numbers of samples ($\mathcal{K}>n$). A more efficient approach derives $\mathcal{Q}_{n}$ as:
$\mathcal{Q}_{n}\equiv\frac{\mathcal{R}_{1}+\mathcal{R}_{2}+\cdots+\mathcal{R}_{n-1}}{n-1}$ (7)
We can express the update rule for action value recursively
$\mathcal{Q}_{n+1}=\mathcal{Q}_{n}+\frac{1}{n}[\mathcal{R}_{n}-\mathcal{Q}_{n}]$ (8)
@@
This recursive equation requires memory only for $\mathcal{Q}_{n}$ and $n$, with minimal computation for each reward [7]. Sample averages apply to stationary bandit problems. In non-stationary environments, recent rewards are more relevant. Using a decaying constant step-size parameter, issues related to non-stationary problems can be mitigated [1].
@@
The initial estimates of action values, $\mathcal{Q}_{1}(a)$, play a crucial role in the learning process. These initial guesses influence the early decisions made by the agent. While sample-average methods can reduce this initial bias after each action is chosen at least once, methods using a constant step-size parameter, $\alpha$, tend to mitigate this bias more gradually over time. Setting optimistic initial values can be particularly advantageous. By assigning higher initial estimates (e.g., +5), the agent is encouraged to explore more actions early on. This is because the initial optimism makes untried actions appear more attractive, thus promoting exploration even when the agent uses a greedy strategy. This approach helps ensure that the agent thoroughly investigates the action space before converging to a final policy. However, this strategy requires careful consideration in defining the initial values, which are often set to zero in standard practice. The choice of initial values should reflect an informed guess about the potential rewards, and overly optimistic values can prevent the agent from converging efficiently if not properly managed. Overall, optimistic initial values can be a useful technique to balance exploration and exploitation in RL, encouraging broader exploration and potentially leading to more optimal long-term policies [8].
Over the next subsection, another fundamental concept in RL, MDPs, are introduced.
2.2 Markov Decision Process (MDPs)